a rectangle has a length that is twice its width. the perimeter of the rectangle is 420. find the length of the diagonal of the rectangle
Answer by ewatrrr(24785)


Hi,

Hi,
Let x AND 2x represent the width and length respecively
Question states;
2(2x) + 2x = 420
4x + 2x = 420
6x = 420
x = 70, the width. Length is 140 (2*70)
Applying the Pythagorean theorem
70^2 + 140^2 = d^2
156.5248 is the length of the diagonal